The Amherst Company has a net profits of $11 million, sales of $194 million, and 4.8 million shares of common stock outstanding. The company has total assets of $115 million and total stockholders equity of $54 mlion. It pays $1.08 per share in common dividends, and the stock trades at $20 per share. Given this information, determine the following: a. Amhersts EPS b. Amhersts book value per share and price-to-book-value ratio c. The firms P/E ratio. d. The companys net profit margin e. The stocks dividend payout ratio and its dividend yield. f. The stocks PEG ratio, given that the companys earnings have been growing at an average annual rate of 8.8%
Answer #1

Answer a.

Earnings per share = Net Profits / Number of stock outstanding
Earnings per share = $11.00 million / 4.80 million
Earnings per share = $2.29

Answer b.

Book Value per share = Total Stockholders’ Equity / Number of stock outstanding
Book Value per share = $54.00 million / 4.80 million
Book Value per share = $11.25

Price-to-book-value Ratio = Market Price per share / Book Value per share
Price-to-book-value Ratio = $20.00 / $11.25
Price-to-book-value Ratio = 1.78

Answer c.

P/E Ratio = Market Price per share / Earnings per share
P/E Ratio = $20.00 / $2.29
P/E Ratio = 8.73

Answer d.

Net Profit Margin = Net Profits / Sales
Net Profit Margin = $11.00 million / $194 million
Net Profit Margin = 5.67%

Answer e.

Dividend Payout Ratio = Dividend per share / Earnings per share
Dividend Payout Ratio = $1.08 / $2.29
Dividend Payout Ratio = 47.16%

Dividend Yield = Dividend per share / Market Price per share
Dividend Yield = $1.08 / $20.00
Dividend Yield = 5.40%

Answer f.

PEG Ratio = P/E Ratio / Growth Rate
PEG Ratio = 8.73 / 8.80
PEG Ratio = 0.99